# Gait Transitions in Load-Pulling Quadrupeds: Insights from Sled Dogs and a Minimal SLIP Model

### v1

@@ II.2 Gait Pattern Definitions
[…]
@@
Phase lags of zero correspond to bounding or half-bounding gaits, while values of $0.5$ indicate symmetrical gaits such as pacing or trotting, following Hildebrand’s classification [4]. Within this space, each canonical galloping gait $G_{e}$, for $e\in\{\textit{TL},\textit{TR},\textit{RL},\textit{RR}\}$, is uniquely determined by the sign pair $(d_{f},d_{h})$:
  • TL (transverse, left-leading): $d_{f}>0,\ d_{h}>0$
  • TR (transverse, right-leading): $d_{f}<0,\ d_{h}<0$
  • RL (rotary, left-leading): $d_{f}>0,\ d_{h}<0$
  • RR (rotary, right-leading): $d_{f}<0,\ d_{h}>0$
These gait labels are visually depicted in Fig. 2. Transverse gallops appear in the top row (white background), and rotary gallops in the bottom row (grey background). Colored bars indicate stance phases; schematic illustrations below each plot highlight limb movement; and dot plots emphasize the sequence of footfalls and leading limb. Gait transitions correspond to sign changes in either $d_{f}$ or $d_{h}$. For instance, a transition from TL to RL involves a reversal in hindlimb phasing (change in $d_{h}$), while a switch from RL to RR reflects a forelimb phase reversal (change in $d_{f}$). Up to 12 distinct transitions can be theoretically defined among the four canonical galloping gaits based on the sign combinations of phase lags. These transitions form the basis for analyzing discrete gait-switching behaviors observed in sled dogs.
@@ III.1 Quadrupedal SLIP Model with Load
To enable gait transitions and capture limb-specific asymmetries observed in natural locomotion, each leg is assigned an independent swing-leg stiffness parameter $k_{\rm swing}$, which may vary between consecutive strides. For an $N$-stride sequence, the stiffness set is denoted as $\bm{k}_{\rm swing}=\{k_{\mathrm{swing},i}^{n}\}$, where $i\in\{\small\text{LH},\small\text{LF},\small\text{RF},\small\text{RH}\}$ and $n=1{:}N$. This formulation supports stride-wise modulation of limb stiffness and enables replication of experimentally observed gait transitions driven by localized changes in limb dynamics.
@@ III.2 Hybrid Dynamics of the Quad-load System
We assume that each leg undergoes exactly one touchdown and one liftoff per stride to avoid Zeno behavior [30, 31]. As a result, the eight timing variables $t_{i}$ in (1) partition the hybrid gait cycle into domains $j=1{:}9$, each corresponding to a unique combination of ground-contacting legs. Within each domain, the associated holonomic constraints are enforced using stacked Jacobians $\bm{J}_{j}$ [18], yielding the following differential-algebraic formulation:
@@
where $\bm{\lambda}_{j}$ includes all GRFs in domain $\mathcal{F}_{j}$. Each of the $162^{4}=$ 16 possible contact combinations defines a continuous domain in the hybrid system.
@@
:$\mathcal{H}:\left\{\begin{array}[]{ll}\mathcal{F}_{j},&(\bm{q},\dot{\bm{q}})\notin\mathcal{C}_{j}(t),\\ \dot{\bm{q}}^{+}=\Delta_{\mathcal{F}_{j}\rightarrow\mathcal{F}_{j+1}}\dot{\bm{q}}^{-},&(\bm{q},\dot{\bm{q}})\in\mathcal{C}_{j}(t),\end{array}\right.$ (6)
@@ III.3 Trajectory Optimization for Gait Replication
$\min_{\begin{subarray}{c}\bm{q}_{0},\ \dot{\bm{q}}_{0},\ \bm{p}\\ \end{subarray}}$ $\ \ \sum_{n=1}^{N}\Big{[}w_{d}\left\|T_{\mathrm{sim}}^{n}-T_{\mathrm{exp}}^{n}\right\|^{2}+w_{t}\left\|\bm{t}_{\mathrm{sim}}^{n}-\bm{t}_{\mathrm{exp}}^{n}\right\|^{2}$
$\ \ +w_{f}\left\|\bm{F}_{\mathrm{sim}}^{n}-\bm{F}_{\mathrm{exp}}^{n}\right\|^{2}+w_{r}\left\|\bm{R}^{n}(\bm{q})\right\|^{2}\Big{]},$ (7)
@@
Multi-Stride Optimization: For gait transitions and sequence replication, the trajectory is segmented into $N$ strides indexed by $n=1{:}N$, with each stride spanning $t_{0}^{n}$ to $t_{\mathrm{end}}^{n}$. Continuity is enforced between successive strides:
@@ IV.1 Observed Distribution of Gait Patterns
Our analysis of high-speed video recordings revealed that sled dogs frequently switch between different galloping gaits during load-pulling. Fig. 4 summarizes the observed distribution of gallops and the frequency of transitions between them for two individual dogs. The arrangement of gait types in the figure follows the same convention as Fig. 2, with transverse gaits (TL, TR) placed above and rotary gaits (RL, RR) below, and left- and right-leading gaits positioned to the left and right, respectively. The size of each dot is proportional to the number of strides recorded for that gait, while the thickness of the connecting arrows reflects the frequency of transitions between gaits.
@@
Data for Individual 1, compiled from 189 recorded strides, reveals the use of all four galloping gait patterns. Among these, RL was the most frequently used, occurring in 120 out of 190 strides. This was followed by TR, which appeared in 50 strides. In contrast, TL and RR were used only occasionally. Gait transitions were unevenly distributed. For RL, 79.83% of consecutive strides remained in the same gait. Transitions to TR occurred in 15.13% of cases, and to TL in 5.04%. For strides beginning in TR, 44.00% remained in TR, 50.00% transitioned to RL, and only 6.00% transitioned to TL. No transitions to RR were observed from TR either. These statistics suggest a strong preference for RL and TR, while TL and RR appear to be infrequent or short-lived. The most common transitions occurred between RL and TR (RL $\leftrightarrow$ TR), which require changing the touchdown order of only the hindlimbs. In contrast, transitions between RL and RR were never observed, likely because they involve a simultaneous reordering of both the forelimb and hindlimb touchdown sequences within a single stride. This more complex coordination change may be less favorable during steady high-speed locomotion.
@@
Data for Individual 2, compiled from 195 recorded strides, reveals a highly consistent gait preference. Nearly all strides were executed in the TR gait, with only a small fraction performed in RL. Specifically, TR accounted for 209 out of 211 strides, while RL was observed in just a single stride. The other two gaits, TL and RR, were not used at all in the recorded dataset. No transitions were observed among three of the four gait types, as TR and RL were the only patterns present. For strides beginning in TR, 99.52% remained in the same gait, while only 0.48%, corresponding to a single stride, transitioned to RL. Since RL occurred only once and was not followed by any additional strides of that type, no outbound transitions from RL could be evaluated. This consistency indicates a strong and stable preference for the TR gait in Individual 2, with other patterns either avoided or unsustainable. The lack of bidirectional transitions suggests that, in contrast to Individual 1, this dog maintained a highly stable coordination strategy with minimal gait variability across the recorded session.
@@ IV.2 Single Stride Replication
In this subsection, we applied the optimization algorithm described in Section III.3 to replicate the four galloping gaits observed in our dataset. We compared footfall timings and tugline forces for both Individual 1 and Individual 2 across the gaits RL, RR, TL, and TR.
@@
For Individual 1, the two most frequently used gaits, RL and TR, are shown in Fig. 5 for detailed comparison between empirical observations and model simulation. Mean touchdown and liftoff timings were extracted from the recorded stride sequences. The averaged stance durations for each leg are shown as white bars in the left panels of Fig. 5, with error bars indicating variability. Simulated stance durations are overlaid as colored bars for direct comparison. In the right panels, tugline forces are shown as solid blue curves for the simulation and as dashed red curves with shaded bands indicating the measured mean and variance.
@@
Figure 5: Single-stride solutions for Individual 1, comparing two gaits: (A) RL and (B) TR. Left subplots show experimental and simulated footfall sequences; right subplots illustrate experimental and simulated tugline forces along the tugline. Experimental means (solid curves) and variances (shaded regions and error bars) are computed from 120 (RL) and 50 (TR) samples.
@@
Table 2: Summary of $R^{2}$ evaluations for footfall timing ($R^{2}_{t}$), tugline force ($R^{2}_{F}$), and stride duration ($R^{2}_{D}$) across different gait types and individuals. The number of strides used to compute each experimental average is listed in the rightmost column. Data for RL and TR for Individual 1 correspond to those shown in Fig. 5.
Dog Name Gait Type $R^{2}_{t}$ $R^{2}_{F}$ $R^{2}_{D}$ Strides
Individual 1 TL 0.99 0.95 0.93 9
Individual 1 TR 0.99 0.96 0.99 50
Individual 1 RL 0.99 0.96 1.00 120
Individual 1 RR 0.99 0.95 0.95 10
Individual 2 TR 0.99 0.95 0.92 209
Individual 2 RL 0.99 0.96 0.99 2
@@
For the RL gait of Individual 1 (Fig. 5A), the simulated stance onset and offset closely match the experimental profiles across all limbs. The mean stance duration error is below 3% of the stride cycle for each leg. The tugline force profile also aligns well, with the simulated curve capturing both the timing and amplitude of the force peaks. The peak force in simulation reaches approximately $0.27\,mg$, consistent with the experimental average. Minor discrepancies appear in the force rising phase but remain within experimental variability. This high degree of agreement is reflected in Table 2, where RL yields $R^{2}=0.99$ for footfall timing, $R^{2}=0.96$ for tugline force, and a perfect $R^{2}=1.00$ for stride duration.
@@
As for the TR gait of Individual 1 (Fig. 5B), the model again reproduces stance durations with high accuracy. Some asymmetry in the forelimb stance timing is evident in the experimental data and is qualitatively captured by the simulation. Tugline force agreement is similarly strong, with an overall $R^{2}=0.96$ and peak force amplitude closely matching the measured mean. The stride duration agreement is slightly lower than for RL but still excellent, at $R^{2}=0.99$.
@@
Due to space constraints, the optimization results for the remaining gaits and for Individual 2 are not visualized here but are included in the publicly available source code and data repository . Table 2 summarizes the $R^{2}$ values across all gait and individual combinations. Footfall timing predictions are highly consistent, with all combinations achieving $R^{2}_{t}=0.99$, indicating precise alignment of simulated and experimental touchdown and liftoff events. Tugline force predictions also show strong agreement, with $R^{2}_{F}$ values ranging from 0.95 to 0.96 across gaits. Stride duration tracking is most accurate for the frequently used RL and TR gaits, where $R^{2}_{D}$ exceeds 0.99, and is slightly lower for TL and RR, which exhibit more variability. Although TL and RR are less frequently used gaits, the SLIP model still maintains reasonable fidelity, with timing and force $R^{2}$ values above 0.93 in all cases. Slightly reduced agreement in TL, especially for Individual 1, may reflect increased variability in experimental execution or decreased model robustness for rarely used gaits. Nevertheless, the model performs reliably across all conditions tested.
@@
Figure 6: Replicated gait transition for Individual 1 from transverse right (TR) to rotary left (RL) gallop. (A) Simulated footfall sequence. (B) Swing leg angular velocities. (C) Tugline force trajectory. (D) Absolute cost variation under ±$\pm 5\%$ perturbations in $\Delta k_{\text{swing},i}$. (E) Relative percentage cost variation. These results highlight the dominant role of hindlimb stiffness modulation in driving the transition.
@@ IV.3 Gait Transition Replication
TR $\leftrightarrow$ RL Gait Transition
As shown in Fig. 6(A), the simulation accurately reproduces the key mechanical features of the TR $\rightarrow$ RL transition, achieving a weighted coefficient of determination of $R^{2}_{\text{w}}=0.95$. Both the footfall sequence and the tugline force trajectory closely match experimental observations across the two-stride transition. The middle panel illustrates swing leg angular velocities, which exhibit near-sinusoidal trajectories during flight—consistent with passive leg oscillations. Since the oscillation period inversely reflects swing stiffness, shorter cycles indicate higher values of $k_{\text{swing}}$.
@@
To evaluate whether the identified stiffness pattern is both necessary and sufficient for achieving the transition, we performed a sensitivity analysis by perturbing each $\Delta k_{\text{swing}}$ by ±$\pm 5\%$ and observing changes in the objective function. As shown in Fig. 6(B, top), the hindlimb curves (LH and RH) show clear minima at zero perturbation, indicating local optimality. In contrast, the forelimb curves (LF and RF) exhibit much smaller variations. The maximum cost increase was 0.54% for LH and 0.39% for RH, compared to only 0.13% and 0.06% for LF and RF, respectively. These results confirm that the transition is highly sensitive to hindlimb stiffness modulation, while the forelimbs remain largely uninvolved.
@@
To further test the sufficiency of this mechanism, we simulated the reverse transition (RL $\rightarrow$ TR). The resulting stiffness changes were $+12.5\,mgl_{0}$ for LH and $-17.7\,mgl_{0}$ for RH, again showing large, opposing adjustments in the hindlimbs. The forelimb changes remained small: $-4.1\,mgl_{0}$ for LF and $+0.72\,mgl_{0}$ for RF. These results mirror the asymmetry observed in the forward transition and reinforce the conclusion that abrupt hindlimb stiffness modulation is sufficient to switch footfall sequences. The full set of values is provided in Table 3, using the same highlighting scheme.
@@
Together, these results show that the TR $\leftrightarrow$ RL transition can be achieved through hindlimb stiffness modulation alone, without requiring adjustments in the forelimbs. This supports the broader hypothesis that discrete gait transitions in quadrupeds can be driven by targeted, stride-scale mechanical adjustments in a small subset of limbs.
RR $\leftrightarrow$ TR Gait Transition
@@
Figure 7: Replicated gait transition for Individual 1 from rotary right (RR) to transverse right (TR) gallop. (A) Simulated footfall sequence. (B) Swing leg angular velocities. (C) Tugline force trajectory. (D) Absolute cost variation under ±$\pm 5\%$ perturbations in $\Delta k_{\text{swing},i}$. (E) Relative percentage cost variation. These results highlight the dominant role of forelimb stiffness modulation in enabling the transition.
@@
Figure 7(A) illustrates the simulated dynamics for the RR $\rightarrow$ TR transition. The model achieves excellent agreement with experimental footfall and tugline force data, yielding a weighted fit score of $R^{2}_{\text{w}}=0.99$. In the middle panel, the right fore (RF, yellow) initially oscillates faster than the left fore (LF, red), and this asymmetry increases in the second stride—RF becomes even faster, while LF slows. In contrast, the hindlimb angular velocities remain nearly identical across both strides, suggesting that the transition is primarily driven by forelimb stiffness modulation.
@@
This interpretation is supported by the corresponding changes in swing stiffness (Table 3). The forelimbs exhibit strong, opposing changes: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LF}}=+9.5\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RF}}=-15.2\,mgl_{0}$. The hindlimb changes are minimal: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LH}}=-1.8\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RH}}=-0.5\,mgl_{0}$. This pattern, highlighted in the table with shaded forelimb cells, contrasts sharply with the hindlimb-driven transition observed in the TR $\leftrightarrow$ RL case. Sensitivity analysis (Fig. 7(B)) confirms the dominant role of the forelimbs. Perturbations of $\Delta k_{\mathrm{swing},\ i}$ in LF and RF led to cost variations of 0.65% and 0.41%, respectively, compared to just 0.13% and 0.08% for LH and RH. These findings demonstrate that precise modulation of forelimb swing stiffness is sufficient to reproduce the observed footfall sequence transition.
@@
Since the reverse transition (TR $\rightarrow$ RR) was not observed in our dataset, we limited our analysis to the RR $\rightarrow$ TR case. Nonetheless, the consistency of the stiffness modulation and the high reconstruction fidelity reinforce our broader hypothesis that discrete gait transitions can be achieved through stride-scale adjustments to a subset of limb parameters, in this case via targeted forelimb stiffness adaptation.

### v2

@@ II.2 Gait Pattern Definitions
$\bm{t}=[t^{\scriptstyle\text{TD}}_{\scriptstyle\text{LH}},\ t^{\scriptstyle\text{LO}}_{\scriptstyle\text{LH}},\ t^{\scriptstyle\text{TD}}_{\scriptstyle\text{LF}},\ t^{\scriptstyle\text{LO}}_{\scriptstyle\text{LF}},\ t^{\scriptstyle\text{TD}}_{\scriptstyle\text{RF}},\ t^{\scriptstyle\text{LO}}_{\scriptstyle\text{RF}},\ t^{\scriptstyle\text{TD}}_{\scriptstyle\text{RH}},\ t^{\scriptstyle\text{LO}}_{\scriptstyle\text{RH}}].$ (1)
@@
Phase lags of zero correspond to bounding or half-bounding gaits, while values of $0.5$ indicate symmetrical gaits such as pacing or trotting, following Hildebrand’s classification [4]. Within this space, each canonical galloping gait $G_{e}$, for $e\in\{\textit{TranL},\textit{TranR},\textit{RotL},\textit{RotR}\}$, is uniquely determined by the sign pair $(d_{f},d_{h})$:
  • TranL (transverse, left-leading): $d_{f}>0,\ d_{h}>0$
  • TranR (transverse, right-leading): $d_{f}<0,\ d_{h}<0$
  • RotL (rotary, left-leading): $d_{f}>0,\ d_{h}<0$
  • RotR (rotary, right-leading): $d_{f}<0,\ d_{h}>0$
These gait labels are visually depicted in Fig. 2. Transverse gallops appear in the top row (white background), and rotary gallops in the bottom row (grey background). Colored bars indicate stance phases; schematic illustrations below each plot highlight limb movement; and dot plots emphasize the sequence of footfalls and leading limb. Gait transitions correspond to sign changes in either $d_{f}$ or $d_{h}$. For instance, a transition from TranL to RotL involves a reversal in hindlimb phasing (change in $d_{h}$), while a switch from RotL to RotR reflects a forelimb phase reversal (change in $d_{f}$). Up to 12 distinct transitions can be theoretically defined among the four canonical galloping gaits based on the sign combinations of phase lags. These transitions form the basis for analyzing discrete gait-switching behaviors observed in sled dogs.
@@ III.1 Quadrupedal SLIP Model with Load
To enable gait transitions and capture limb-specific asymmetries observed in natural locomotion, each leg is assigned an independent swing-leg stiffness parameter $k_{\rm swing}$, which may vary between consecutive strides. For an $N$-stride sequence, the stiffness set is denoted as $\bm{k}_{\rm swing}=\{k_{\mathrm{swing},i}^{n}\}$, where $i\in\{\small\text{LH},\small\text{LF},\small\text{RF},\small\text{RH}\}$ and n=1:Nn=1{:}N. This formulation supports stride-wise modulation of limb stiffness and enables replication of experimentally observed gait transitions driven by localized changes in limb dynamics.
@@ III.2 Hybrid Dynamics of the Quad-load System
We assume that each leg undergoes exactly one touchdown and one liftoff per stride to avoid Zeno behavior [30, 31]. As a result, the eight timing variables $t_{i}$ in (1) partition the hybrid gait cycle into domains j=1:9j=1{:}9, each corresponding to a unique combination of ground-contacting legs. Within each domain, the associated holonomic constraints are enforced using stacked Jacobians $\bm{J}_{j}$ [18], yielding the following differential-algebraic formulation:
@@
where $\bm{\lambda}_{j}$ includes all GRFs in domain $\mathcal{F}_{j}$. Each of the $2^{4}=16$ possible contact combinations defines a continuous domain in the hybrid system.
@@
$\mathcal{H}:\left\{\begin{array}[]{ll}\mathcal{F}_{j},&(\bm{q},\dot{\bm{q}})\notin\mathcal{C}_{j}(t),\\ \dot{\bm{q}}^{+}=\Delta_{\mathcal{F}_{j}\rightarrow\mathcal{F}_{j+1}}\dot{\bm{q}}^{-},&(\bm{q},\dot{\bm{q}})\in\mathcal{C}_{j}(t),\end{array}\right$. (6)
@@ III.3 Trajectory Optimization for Gait Replication
$\min_{\begin{subarray}{c}\bm{q}_{0},\ \dot{\bm{q}}_{0},\ \bm{p}\\ \end{subarray}}$ $\ \ \sum_{n=1}^{N}\Big{[}w_{t}\left\|\bm{t}_{\mathrm{sim}}^{n}-\bm{t}_{\mathrm{exp}}^{n}\right\|^{2}+w_{f}\left\|\bm{F}_{\mathrm{sim}}^{n}-\bm{F}_{\mathrm{exp}}^{n}\right\|^{2}$
$\ \ +w_{d}\left\|T_{\mathrm{sim}}^{n}-T_{\mathrm{exp}}^{n}\right\|^{2}+w_{r}\left\|\bm{R}^{n}(\bm{q})\right\|^{2}\Big{]},$ (7)
@@
Multi-Stride Optimization: For gait transitions and sequence replication, the trajectory is segmented into $N$ strides indexed by n=1:Nn=1{:}N, with each stride spanning $t_{0}^{n}$ to $t_{\mathrm{end}}^{n}$. Continuity is enforced between successive strides:
@@ IV.1 Observed Distribution of Gait Patterns
Our analysis of high-speed video recordings revealed that sled dogs frequently switch between different galloping gaits during load-pulling. Fig. 4 summarizes the observed distribution of gallops and the frequency of transitions between them for two individual dogs. The arrangement of gait types in the figure follows the same convention as Fig. 2, with transverse gaits (TranL, TranR) placed above and rotary gaits (RotL, RotR) below, and left- and right-leading gaits positioned to the left and right, respectively. The size of each dot is proportional to the number of strides recorded for that gait, while the thickness of the connecting arrows reflects the frequency of transitions between gaits.
@@
Data for Individual 1, compiled from 189 recorded strides, reveals the use of all four galloping gait patterns. Among these, RotL was the most frequently used, occurring in 120 out of 190 strides. This was followed by TranR, which appeared in 50 strides. In contrast, TranL and RotR were used only occasionally. Gait transitions were unevenly distributed. For RotL, 79.83% of consecutive strides remained in the same gait. Transitions to TranR occurred in 15.13% of cases, and to TranL in 5.04%. For strides beginning in TranR, 44.00% remained in TranR, 50.00% transitioned to RotL, and only 6.00% transitioned to TranL. No transitions to RotR were observed from TranR either. These statistics suggest a strong preference for RotL and TranR, while TranL and RotR appear to be infrequent or short-lived. The most common transitions occurred between RotL and TranR (RotL $\leftrightarrow$ TranR), which require changing the touchdown order of only the hindlimbs. In contrast, transitions between RotL and RotR were never observed, likely because they involve a simultaneous reordering of both the forelimb and hindlimb touchdown sequences within a single stride. This more complex coordination change may be less favorable during steady high-speed locomotion.
@@
Data for Individual 2, compiled from 195 recorded strides, reveals a highly consistent gait preference. Nearly all strides were executed in the TranR gait, with only a small fraction performed in RotL. Specifically, TranR accounted for 209 out of 211 strides, while RotL was observed in just a single stride. The other two gaits, TranL and RotR, were not used at all in the recorded dataset. No transitions were observed among three of the four gait types, as TranR and RotL were the only patterns present. For strides beginning in TranR, 99.52% remained in the same gait, while only 0.48%, corresponding to a single stride, transitioned to RotL. Since RotL occurred only once and was not followed by any additional strides of that type, no outbound transitions from RotL could be evaluated. This consistency indicates a strong and stable preference for the TranR gait in Individual 2, with other patterns either avoided or unsustainable. The lack of bidirectional transitions suggests that, in contrast to Individual 1, this dog maintained a highly stable coordination strategy with minimal gait variability across the recorded session.
@@ IV.2 Single Stride Replication
In this subsection, we applied the optimization algorithm described in Section III.3 to replicate the four galloping gaits observed in our dataset. We compared footfall timings and tugline forces for both Individual 1 and Individual 2 across the gaits RotL, RotR, TranL, and TranR.
@@
For Individual 1, the two most frequently used gaits, RotL and TranR, are shown in Fig. 5 for detailed comparison between empirical observations and model simulation. Mean touchdown and liftoff timings were extracted from the recorded stride sequences. The averaged stance durations for each leg are shown as white bars in the left panels of Fig. 5, with error bars indicating variability. Simulated stance durations are overlaid as colored bars for direct comparison. In the right panels, tugline forces are shown as solid blue curves for the simulation and as dashed red curves with shaded bands indicating the measured mean and variance.
@@
Figure 5: Single-stride solutions for Individual 1, comparing two gaits: (A) TranR and (B) RotL. Left subplots show experimental and simulated footfall sequences; right subplots illustrate experimental and simulated tugline forces along the tugline. Experimental means (solid curves) and variances (shaded regions and error bars) are computed from 50 (TranR) and 120 (RotL) samples.
@@
Table 2: Summary of $R^{2}$ evaluations for footfall timing ($R^{2}_{t}$), tugline force ($R^{2}_{F}$), and stride duration ($R^{2}_{D}$) across different gait types and individuals. The number of strides used to compute each experimental average is listed in the rightmost column. Data for RotL and TranR for Individual 1 correspond to those shown in Fig. 5.
Dog Name Gait Type $R^{2}_{T}$ $R^{2}_{F}$ $R^{2}_{D}$ # Strides
Individual 1 TranL 0.99 0.95 0.89 9
Individual 1 TranR 0.99 0.96 0.99 50
Individual 1 RotL 0.99 0.94 0.93 120
Individual 1 RotR 0.99 0.96 0.99 10
Individual 2 TranR 0.98 0.96 0.84 209
Individual 2 RotL 0.98 0.92 0.87 2
@@
For the RotL gait of Individual 1 (Fig. 5A), the simulated stance onset and offset closely match the experimental profiles across all limbs, indicating precise replication of footfall timing. Quantitatively, the mean stance duration error remains below 3% of the stride cycle for each leg. The tugline force profile also demonstrates strong agreement: the simulated curve captures both the timing and amplitude of the experimental force peaks, with a peak value reaching approximately 0.27mg0.27\,mg, consistent with the observed average. Although minor discrepancies occur during the force rising phase, they remain within the bounds of experimental variability. This high level of fidelity is reflected in the quantitative metrics summarized in Table 2, where the RotL simulation achieves $R^{2}_{T}=0.99$ for footfall timing, $R^{2}_{F}=0.94$ for tugline force, and $R^{2}_{D}=0.93$ for stride duration.
@@
The model performs similarly well for the TranR gait of Individual 1 (Fig. 5B), further demonstrating its robustness across gait types. Stance durations are accurately reproduced, and notable asymmetries in forelimb stance timing—evident in the experimental data—are qualitatively reflected in the simulation. Tugline force agreement remains strong, with an overall $R^{2}_{F}=0.96$, and the simulated peak force amplitude closely aligns with the experimental mean. In terms of timing accuracy, the footfall sequence achieves an excellent match with $R^{2}_{T}=0.99$. Furthermore, the stride duration agreement is nearly perfect, as indicated by a coefficient of determination of $R^{2}_{D}=0.99$.
@@
Due to space constraints, the optimization results for the remaining gaits and for Individual 2 are not visualized here but are included in the publicly available source code and data repository . Table 2 summarizes the $R^{2}$ values across all gait and individual combinations. Footfall timing predictions are highly consistent, with all combinations achieving $R^{2}_{t}=0.99$, indicating precise alignment of simulated and experimental touchdown and liftoff events. Tugline force predictions also show strong agreement, with $R^{2}_{F}$ values ranging from 0.95 to 0.96 across gaits. Stride duration tracking is most accurate for the frequently used RotL and TranR gaits, where $R^{2}_{D}$ exceeds 0.99, and is slightly lower for TranL and RotR, which exhibit more variability. Although TranL and RotR are less frequently used gaits, the SLIP model still maintains reasonable fidelity, with timing and force $R^{2}$ values above 0.93 in all cases. Slightly reduced agreement in TranL, especially for Individual 1, may reflect increased variability in experimental execution or decreased model robustness for rarely used gaits. Nevertheless, the model performs reliably across all conditions tested.
@@
Figure 6: Replicated gait transition for Individual 1 from transverse right (TranR) to rotary left (RotL) gallop. (A) Simulated footfall sequence. (B) Swing leg angular velocities. (C) Tugline force trajectory. (D) Absolute cost variation under $\pm 5\%$ perturbations in $\Delta k_{\text{swing},i}$. (E) Relative percentage cost variation. These results highlight the dominant role of hindlimb stiffness modulation in driving the transition.
@@ IV.3 Gait Transition Replication
TranR $\leftrightarrow$ RotL Gait Transition
As shown in Fig. 6(A), the simulation accurately reproduces the key mechanical features of the TranR $\rightarrow$ RotL transition, achieving a weighted coefficient of determination of $R^{2}_{\text{w}}=0.85$. Both the footfall sequence and the tugline force trajectory closely match experimental observations across the two-stride transition. The middle panel illustrates swing leg angular velocities, which exhibit near-sinusoidal trajectories during flight—consistent with passive leg oscillations. Since the oscillation period inversely reflects swing stiffness, shorter cycles indicate higher values of $k_{\text{swing}}$.
@@
To evaluate whether the identified stiffness pattern is both necessary and sufficient for achieving the transition, we performed a sensitivity analysis by perturbing each $\Delta k_{\text{swing}}$ by $\pm 5\%$ and observing changes in the objective function. As shown in Fig. 6(B, top), the hindlimb curves (LH and RH) show clear minima at zero perturbation, indicating local optimality. In contrast, the forelimb curves (LF and RF) exhibit much smaller variations. The maximum cost increase was 0.54% for LH and 0.39% for RH, compared to only 0.13% and 0.06% for LF and RF, respectively. These results confirm that the transition is highly sensitive to hindlimb stiffness modulation, while the forelimbs remain largely uninvolved.
@@
To further test the sufficiency of this mechanism, we simulated the reverse transition (RotL $\rightarrow$ TranR). The resulting stiffness changes were $12.5\,mgl_{0}$ for LH and $17.7\,mgl_{0}$ for RH, again showing large, opposing adjustments in the hindlimbs. The forelimb changes remained small: $4.1\,mgl_{0}$ for LF and $0.72\,mgl_{0}$ for RF. These results mirror the asymmetry observed in the forward transition and reinforce the conclusion that abrupt hindlimb stiffness modulation is sufficient to switch footfall sequences. The full set of values is provided in Table 3, using the same highlighting scheme.
@@
Together, these results show that the TranR $\leftrightarrow$ RotL transition can be achieved through hindlimb stiffness modulation alone, without requiring adjustments in the forelimbs. This supports the broader hypothesis that discrete gait transitions in quadrupeds can be driven by targeted, stride-scale mechanical adjustments in a small subset of limbs.
RotL $\leftrightarrow$ TranL Gait Transition
@@
Figure 7: Replicated gait transition for Individual 1 from rotary left (RotL) to transverse left (TranL) gallop. (A) Simulated footfall sequence. (B) Swing leg angular velocities. (C) Tugline force trajectory. (D) Absolute cost variation under $\pm 5\%$ perturbations in $\Delta k_{\text{swing},i}$. (E) Relative percentage cost variation. These results highlight the dominant role of forelimb stiffness modulation in enabling the transition.
@@
Figure 7(A) illustrates the simulated dynamics for the RotL $\rightarrow$ TranL transition. The model achieves excellent agreement with experimental footfall and tugline force data, yielding a weighted fit score of $R^{2}_{\text{w}}=0.96$. In the middle panel, the left fore (LF, red) initially oscillates faster than the right fore (RF, yrllow), and this asymmetry increases in the second stride—RF becomes even faster, while LF slows. In contrast, the hindlimb angular velocities remain nearly identical across both strides, suggesting that the transition is primarily driven by forelimb stiffness modulation.
@@
This interpretation is supported by the corresponding changes in swing stiffness (Table 3). The forelimbs exhibit strong, opposing changes: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LF}}=-9.5\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RF}}=15.2\,mgl_{0}$. The hindlimb changes are minimal: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LH}}=-1.8\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RH}}=-0.5\,mgl_{0}$. This pattern, highlighted in the table with shaded forelimb cells, contrasts sharply with the hindlimb-driven transition observed in the TranR $\leftrightarrow$ RotL case. Sensitivity analysis (Fig. 7(B)) confirms the dominant role of the forelimbs. Perturbations of $\Delta k_{\mathrm{swing},\ i}$ in LF and RF led to cost variations of 0.65% and 0.41%, respectively, compared to just 0.13% and 0.08% for LH and RH. These findings demonstrate that precise modulation of forelimb swing stiffness is sufficient to reproduce the observed footfall sequence transition.
@@
Since the reverse transition (TranL $\rightarrow$ RotL) was not observed in our dataset, we limited our analysis to the RotL $\rightarrow$ TranL case. Nonetheless, the consistency of the stiffness modulation and the high reconstruction fidelity reinforce our broader hypothesis that discrete gait transitions can be achieved through stride-scale adjustments to a subset of limb parameters, in this case via targeted forelimb stiffness adaptation.